*Hand in at the BEGINNING of 1. In 1.0 mole of the following, how many atoms do you have? 1 mol Mg = atoms Mg 1 mol Ag = atoms Ag 1 mol Au = atoms Au 1 mol Pb = atoms Pb 2. In 1.0 mole of the following, how many molecules do you have? 1 mol KCI molecules KCI molecules MgCl2 1 mol MgCl2 = _molecules C12H22O11 (sugar) 1 mol C12H22O11 = 3. If you were to measure out 1.0 mole of each of the following, how many grar have (use 2 decimal places for values)? g Mg 1 mol Mg = .g Ag 1 mol Ag = -8 Au
Answer #1

1)

1 mol of any compound have Avogadro’s number times number of atom

So,

1 mol of Mg = 6.022*10^23 atoms of Mg

1 mol of Ag = 6.022*10^23 atoms of Ag

1 mol of Au = 6.022*10^23 atoms of Au

1 mol of Pb = 6.022*10^23 atoms of Pb

Only 1 question at a time please